# Closed-Form Feedback-Free Learning with Forward Projection

### v3

@@ Abstract
State-of-the-art methods for backpropagation-free learning employ local error feedback to direct iterative optimisation via gradient descent. In this study, we examine the more restrictive setting where retrograde communication from neuronal outputs is unavailable for pre-synaptic weight optimisation. To address this challenge, we propose Forward Projection (FP). This randomised closed-form training method requires only a single forward pass over the entire dataset for model fitting, without retrograde communication. Our method generates target values for pre-activation membrane potentials at each layer through randomised nonlinear projections of pre-synaptic inputs and the labels, thereby encoding information from both sources. Local loss functions are optimised over pre-synaptic inputs using closed-form regression, without feedback from neuronal outputs or downstream layers. Interpretability is a key advantage of FP training; membrane potentials of hidden neurons in FP-trained networks encode information which are interpretable layer-wise as label predictions. We demonstrate the effectiveness of FP across four biomedical datasets, comparing it with backpropagation and local learning techniques such as Forward-Forward training and Local Supervision in multi-layer perceptron and convolutional architectures. In some few-shot learning tasks, FP yielded more generalisable models than those optimised via backpropagation. In large-sample tasks, FP-based models achieve generalisation comparable to gradient descent-based local learning methods while requiring only a single forward propagation step, achieving significant speed up for training. Interpretation functions defined on local neuronal activity in FP-based models successfully identified clinically salient features for diagnosis in two distinct datasets – identification of key signatures of myocardial infarction in time-series data from electrocardiogram sequences and detection of regions of choroid neovascularisation in optical coherence tomography images.
Forward Projection is a computationally efficient machine learning approach that yields interpretable neural network models without retrograde communication of neuronal activity during training.
@@ Introduction
A core task in neural network training is synaptic “credit assignment” for error in downstream layers [1, 2, 3]. Although backpropagation has been established as the standard approach to this problem, its biological plausibility has been questioned [4, 3, 5]. Backpropagation requires bidirectional synaptic communication, which is incompatible with the unidirectional transmission of neural action potentials [3]. Consequently, backpropagation would require either symmetric neural connectivity or a parallel retrograde network for error feedback to earlier layers [6, 3, 7]. The backward pass traverses layers in reverse order of activation, leading to temporal discordance between forward and backward operations and necessitating storage of hidden activations. Furthermore, gradient descent requires hidden neural activations to be differentiable throughout. To approximate backpropagation in a more biologically plausible way, approaches such as Predictive Coding [8, 5] and Difference Target Propagation [1] seek to align forward neural activity with a backward network that mirrors the forward architecture, effectively implementing an inverse model. Various strategies have been proposed to address the credit assignment problem with reduced retrograde communication requirements. Auxiliary loss functions computed on the activations of individual layers have been proposed to shorten the backward pass [9, 10, 11]. A prominent example is the layer-wise greedy optimisation method that aims to minimise the costs of backpropagation applying locally supervision (LS) at each layer [9, 12]. Extensions of this approach include deep continuous local learning [10] and single layer updating [13]. Hebbian learning rules using asymmetric feedback weights have been explored, allowing for independent updates for forward and feedback pathways [14].
@@
Recent approaches have also explored using two forward passes to facilitate communication between upstream and downstream neurons [15, 16, 17, 18]. The “Forward-Forward” (FF) learning algorithm [15], is an approach in which data and label hypotheses are combined as inputs, with optimisation seeking to upregulate neural response to correctly labelled inputs and subdue responses to spuriously labelled inputs. However, inference under the original forward learning algorithm requires a forward pass for each hypothesised label, presenting issues for tasks with large label spaces [18]. The “PEPITA” algorithm proposed a preliminary forward pass to predict a label, which is used to generate spurious data-label instances for training [16].
@@
The central issue with local learning methods in deep neural networks is that the optimal activity of hidden neurons is unknown during training, preventing direct observation of local error. Pre-defining activity targets for hidden neurons allows for heuristic local optimization. However, target definition strategies are a topic of ongoing research [19]. Techniques such as local-representation alignment and target propagation introduce target values for hidden activations using limited retrograde communication from downstream neurons [1, 20, 21, 11, 19]. Alternatively, target activities can be set as fixed random label projections computed during the forward pass, thereby permitting direct measurement of the error during the forward pass [11]. However, this approach may lead to highly correlated neuronal activity – a problem known as informational collapse [12]. Although additive noise permits maximal decorrelation of target potentials [1], it is uninformative with respect to the label, potentially impeding model fitting. Approaches such as random neural network features [22] forgo optimisation of hidden nodes entirely, projecting inputs to a random high-dimensional non-linear feature space. However, random feature layers require exponential scaling with respect to the input dimension to support downstream learning [23].
@@
The “Predictive Coding” paradigm provides an insightful perspective on the collective functionality of intermediate neurons, proposing that local neural learning processes optimise the prediction of pre-synaptic neural inputs, minimising the “surprise” of out-of-distribution stimuli [5]. Approaches such as the “Difference Target Propagation” [1] reframe network layers as a series of autoencoders, where intermediate activations represent a series of encodings, transitioning from input information to label information. The issue of informational collapse in locally supervised SGD-trained models has been addressed by optimising the retention of information on the pre-synaptic activity [12]. “Prospective configuration” of target activities reformulates learning under the presumption that idealised adjustments to neuronal activity should be generated via energy minimisation before synaptic adjustment [19]. Synaptic weights may be modified to realise the preconceived activities in response to the given input stimuli [19], thereby transitioning from input information to label information through model layers. “Local neural synchronisation” generates target neural activity for hidden layers by projecting neuronal activity onto periodic basis vectors representative of class labels [24].
@@
Going beyond these approaches, we propose the use of random nonlinear projections of both pre-synaptic inputs and target labels to generate local target activities in the forward pass. The objective of this approach is to develop a neural network training algorithm which requires no retrograde communication. Closed-form regression techniques are applicable in this setting, permitting single-step layer weight computation without error feedback. Thus, weight parameters are determined without backward communication from neuronal outputs or downstream layers. Further advantages of this method include the direct interpretability of hidden neurons with respect to local label predictions and stability in the few-shot setting.
@@ Results
We consider a dataset $\mathcal{D}=\{(\textbf{x}_{i},\textbf{y}_{i})\}_{i=1}^{N}$ from the joint distribution $(X,Y)$. The task is to learn a feed-forward neural network function mapping $X\rightarrow Y$. The model has $L$ layers, with dimensions $m_{0},\ldots m_{L}$, and activations $\textbf{a}_{0},\ldots,\textbf{a}_{L}$, where $\textbf{a}_{0}=\textbf{x}$ and $\textbf{a}_{L}=\textbf{y}$. Each layer is equipped with weights $\textbf{W}_{l}$ to generate membrane potential $\textbf{z}_{l}=\textbf{a}_{l-1}\textbf{W}_{l}$, and activation function $f_{l}:\mathbb{R}\rightarrow\mathbb{R}$ to generate neuronal outputs $\textbf{a}_{l}=f_{l}(\textbf{z}_{l})$. The model prediction is defined as $\hat{\textbf{y}}=f_{L}(\textbf{W}_{L}(\ldots f_{1}(\textbf{W}_{1}\textbf{x}))))$.
@@
To generate target membrane potentials for hidden neurons, we present the Forward Projection (FP) algorithm (Figure 1-A). We propose to combine pre-synaptic inputs and labels using random non-linear projections to generate targets (Figure 1-B). For each training sample, the target potential $\tilde{\textbf{z}}_{l}\in\mathbb{R}^{1\times m_{l}}$ is generated from pre-synaptic inputs $\textbf{a}_{l-1}\in\mathbb{R}^{1\times m_{l-1}}$ and labels $\textbf{y}\in\mathbb{R}^{1\times m_{L}}$ using fixed random projection matrices $\textbf{Q}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ and $\textbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$ such that
$\tilde{\textbf{z}}_{l}=g_{l}(\textbf{a}_{l-1}\textbf{Q}_{l})+g_{l}(\textbf{y}\textbf{U}_{l}),$ (1)
@@
where $g_{l}:\mathbb{R}\rightarrow\mathbb{R}$ is an element-wise non-linear transformation. It is noted here that $g_{l}$ is not necessarily equal to $f_{l}$, the neural activation function. $\textbf{Q}_{l}$ and $\textbf{U}_{l}$ are fixed linear projections drawn from random Gaussian distributions, which are pre-defined before training. The target potential $\tilde{\textbf{z}}_{l}$ allows for a local auxiliary loss function to be defined with respect to the actual membrane potential, $\textbf{z}_{l}$, realised during the forward pass. The randomised combination of pre-synaptic inputs with the target label is inspired by the high-dimensional computing paradigm [25], where fixed random projections are employed to encode information from multiple vector inputs, approximately preserving relative distances according to the Johnson-Lindenstrauss lemma [25]. Accordingly, it is noted that $\tilde{\textbf{z}}_{l}$ approximately encodes both $\textbf{a}_{l-1}$ and y (see remark A.7 in SI). By generating target potentials in a forward manner, the mix of information encoded in each layer is expected to transition incrementally from predominantly input information in early layers to label information in later layers. We propose to define a local loss function $\mathcal{L}_{l}:=\lVert\textbf{Z}_{l}-\tilde{\textbf{Z}_{l}}\rVert$, which may be employed as an objective to optimise $\textbf{W}_{l}$. Synaptic weights for each layer can be computed in a closed-form forward manner, using ridge regression (Figure 1-C), such that
$\textbf{W}_{l}:=(\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}+\lambda\textbf{I})^{-1}(\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}).$ (2)
@@
Here, $\textbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$ and $\tilde{\textbf{Z}}_{l}\in\mathbb{R}^{N\times m_{l}}$ are matrices of pre-synaptic activities and target potentials, respectively, collected over the $N$ samples in $\mathcal{D}$. $\lambda$ is a regularisation term and I is the identity matrix. Observe that $\textbf{A}_{l-1}^{T}\textbf{A}_{l-1}$ and $\textbf{A}_{l-1}^{T}\tilde{\textbf{Z}}_{l}$ terms in (2) may be computed sequentially over instances $\{\textbf{a}_{1,l-1},\ldots,\textbf{a}_{N,l-1}\}\subset\mathbb{R}^{1\times m_{l-1}}$, and $\{\tilde{\textbf{z}}_{1,l-1},\ldots,\tilde{\textbf{z}}_{N,l-1}\}\subset\mathbb{R}^{1\times m_{l}}$ (see remark A.6 in SI). Therefore, memory requirements are independent of $N$, depending only on $m_{l-1}$ and $m_{l}$. The approach of encoding information on both inputs and labels in the hidden layers is inspired by the Predictive Coding and target propagation paradigms [5, 1, 9]. However, unlike previous approaches, no backward communication is required from neuronal outputs to pre-synaptic inputs to achieve the FP fit. As a consequence of this fitting approach, neural membrane potentials in hidden layers of FP-trained models may be interpreted as label predictions (Figure 1-C).
@@
Figure 1: A: Forward projection algorithm for fitting layer weights $\textbf{W}_{1},\ldots,\textbf{W}_{l}$ to model labels Y from data X. B: Procedure for generating the $l$-th layer target potentials $\tilde{\textbf{Z}}_{l}$. Pre-synaptic inputs $\textbf{A}_{l-1}$ and labels Y are projected with fixed matrices $\textbf{Q}_{l}$ and $\textbf{U}_{l}$, respectively before applying non-linearity $g_{l}$. C: Optimising $\textbf{W}_{l}$ to predict $\tilde{\textbf{Z}}_{l}$ from $\textbf{A}_{l-1}$ by ridge regression with penalty $\lambda$. D: Interpreting membrane potentials $\textbf{z}_{l}$ as a local label prediction $\hat{\textbf{y}}_{l}$ given pre-synaptic inputs $\textbf{a}_{l-1}$ and projection matrices $\textbf{Q}_{l}$ and $\textbf{U}_{l}^{+}$, where $\textbf{U}_{l}^{+}$ is the the pseudo-inverse of $\textbf{U}_{l}$.

### v4

@@ Abstract
State-of-the-art backpropagation-free learning methods employ local error feedback to direct iterative optimisation via gradient descent. Here, we examine the more restrictive setting where retrograde communication from neuronal outputs is unavailable for pre-synaptic weight optimisation. We propose Forward Projection (FP), a randomised closed-form training method requiring only a single forward pass over the dataset without retrograde communication. FP generates target values for pre-activation membrane potentials through randomised nonlinear projections of pre-synaptic inputs and labels. Local loss functions are optimised using closed-form regression without feedback from downstream layers. A key advantage is interpretability: membrane potentials in FP-trained networks encode information interpretable layer-wise as label predictions. Across several biomedical datasets, FP achieves generalisation comparable to gradient descent-based local learning methods while requiring only a single forward propagation step, yielding significant training speedup. In few-shot learning tasks, FP produces more generalisable models than backpropagation-optimised alternatives, with local interpretation functions successfully identifying clinically salient diagnostic features.
@@ Introduction
A core task in neural network training is synaptic “credit assignment” for error in downstream layers [1, 2, 3]. Although backpropagation has been established as the standard approach to this problem, its biological plausibility has been questioned [4, 3, 5]. Backpropagation requires bidirectional synaptic communication, which is incompatible with the unidirectional transmission of neural action potentials [3]. Consequently, backpropagation would require either symmetric neural connectivity or a parallel retrograde network for error feedback to earlier layers [6, 3, 7]. The backward pass traverses layers in reverse order of activation, leading to temporal discordance between forward and backward operations and necessitating storage of hidden activations. Furthermore, gradient descent requires hidden neural activations to be differentiable throughout. To approximate backpropagation in a more biologically plausible way, approaches such as Predictive Coding [8, 5] and Difference Target Propagation [1] seek to align forward neural activity with a backward network that mirrors the forward architecture, effectively implementing an inverse model. Various strategies have been proposed to address the credit assignment problem with reduced retrograde communication requirements. Auxiliary loss functions computed on the activations of individual layers have been proposed to shorten the backward pass [9, 10, 11, 12]. A prominent example is the layer-wise greedy optimisation method that aims to minimise the costs of backpropagation by applying local supervision (LS) at each layer [9, 13]. Extensions of this approach include deep continuous local learning [10] and single layer updating [14]. Hebbian learning rules using asymmetric feedback weights have been explored, allowing for independent updates for forward and feedback pathways [15].
@@
Recent approaches have also explored using two forward passes to facilitate communication between upstream and downstream neurons [16, 17, 18, 19]. The “Forward-Forward” (FF) learning algorithm [16] is an approach in which data and label hypotheses are combined as inputs, with optimisation seeking to upregulate neural response to correctly labelled inputs and subdue responses to spuriously labelled inputs. However, inference under the original Forward-Forward algorithm requires a forward pass for each hypothesised label, presenting issues for tasks with large label spaces [19]. The “PEPITA” algorithm employed a preliminary forward pass to predict a label, which is used to generate spurious data-label instances for training [17].
@@
The central issue with local learning methods in deep neural networks is that the optimal activity of hidden neurons is unknown during training, preventing direct observation of local error. Pre-defining activity targets for hidden neurons allows for heuristic local optimization. However, target definition strategies are a topic of ongoing research [20]. Techniques such as local-representation alignment and target propagation introduce target values for hidden activations using limited retrograde communication from downstream neurons [1, 21, 22, 11, 20]. Alternatively, target activities can be set as fixed random label projections computed during the forward pass, thereby permitting direct measurement of the error during the forward pass [11]. However, this approach may lead to highly correlated neuronal activity – a problem known as informational collapse [13]. Although additive noise permits maximal decorrelation of target potentials [1], it is uninformative with respect to the label, potentially impeding model fitting. Approaches such as random neural network features [23] forgo optimisation of hidden nodes entirely, instead projecting inputs to a random high-dimensional non-linear feature space. However, random feature layers require exponential scaling with respect to the input dimension to support downstream learning [24].
@@
The “Predictive Coding” paradigm provides a perspective on the collective functionality of intermediate neurons, proposing that local neural learning processes optimise the prediction of pre-synaptic neural inputs, minimising the “surprise” of out-of-distribution stimuli [5]. Approaches such as the “Difference Target Propagation” [1] reframe network layers as a series of autoencoders, where intermediate activations represent a series of encodings, transitioning from input information to label information. The issue of informational collapse in locally supervised SGD-trained models has been addressed by optimising the retention of information on the pre-synaptic activity [13]. “Prospective configuration” of target activities reformulates learning under the presumption that idealised adjustments to neuronal activity should be generated via energy minimisation before synaptic adjustment [20]. Synaptic weights may be modified to realise the predetermined activities in response to the given input stimuli [20], thereby transitioning from input information to label information through model layers. “Local neural synchronisation” generates target neural activity for hidden layers by projecting neuronal activity onto periodic basis vectors representative of class labels [25].
@@
Going beyond these approaches, we propose the use of random nonlinear projections of both pre-synaptic inputs and target labels to generate local target activities in the forward pass. The objective of this approach is to develop a neural network training algorithm that requires no retrograde communication. Closed-form regression techniques are applicable in this setting, permitting single-step layer weight computation without error feedback. Thus, weight parameters are determined without backward communication from neuronal outputs or downstream layers. Further advantages of this method include the direct interpretability of hidden neurons with respect to local label predictions and stability in the few-shot setting.
@@ Results
We consider a dataset $\mathcal{D}=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$ from the joint distribution $(X,Y)$. The task is to learn a feed-forward neural network function mapping $X\rightarrow Y$. The model has $L$ layers, with dimensions $m_{0},\ldots,m_{L}$, and activations $\mathbf{a}_{0},\ldots,\mathbf{a}_{L}$, where $\mathbf{a}_{0}=\mathbf{x}$ and $\mathbf{a}_{L}=\mathbf{y}$. Each layer is equipped with weights $\mathbf{W}_{l}$ to generate membrane potentials $\mathbf{z}_{l}=\mathbf{a}_{l-1}\mathbf{W}_{l}$, and activation function $f_{l}:\mathbb{R}\rightarrow\mathbb{R}$ to generate neuronal outputs $\mathbf{a}_{l}=f_{l}(\mathbf{z}_{l})$. The model prediction is defined as $\hat{\mathbf{y}}=f_{L}(\mathbf{W}_{L}(\ldots f_{1}(\mathbf{W}_{1}\mathbf{x}))))$.
@@
To generate target membrane potentials for hidden neurons, we present the Forward Projection (FP) algorithm (Figure 1-A). We propose to combine pre-synaptic inputs and labels using random non-linear projections to generate targets (Figure 1-B). For each training sample, the target potential $\tilde{\mathbf{z}}_{l}\in\mathbb{R}^{1\times m_{l}}$ is generated from pre-synaptic inputs $\mathbf{a}_{l-1}\in\mathbb{R}^{1\times m_{l-1}}$ and labels $\mathbf{y}\in\mathbb{R}^{1\times m_{L}}$ using fixed random projection matrices $\mathbf{Q}_{l}\in\mathbb{R}^{m_{l-1}\times m_{l}}$ and $\mathbf{U}_{l}\in\mathbb{R}^{m_{L}\times m_{l}}$ such that
$\tilde{\mathbf{z}}_{l}=g_{l}(\mathbf{a}_{l-1}\mathbf{Q}_{l})+g_{l}(\mathbf{y}\mathbf{U}_{l}),$ (1)
@@
where $g_{l}:\mathbb{R}\rightarrow\mathbb{R}$ is an element-wise non-linear transformation. It is noted here that $g_{l}$ is not necessarily equal to $f_{l}$, the neural activation function. $\mathbf{Q}_{l}$ and $\mathbf{U}_{l}$ are fixed linear projections drawn from random Gaussian distributions, which are pre-defined before training. The target potential $\tilde{\mathbf{z}}_{l}$ allows a local auxiliary loss function to be defined with respect to the actual membrane potential, $\mathbf{z}_{l}$, realised during the forward pass. The randomised combination of pre-synaptic inputs with the target label is inspired by the high-dimensional computing paradigm [26], where fixed random projections are employed to encode information from multiple vector inputs, approximately preserving relative distances according to the Johnson-Lindenstrauss lemma [26]. Accordingly, it is noted that $\tilde{\mathbf{z}}_{l}$ approximately encodes both $\mathbf{a}_{l-1}$ and $\mathbf{y}$ (see remark A.7 in SI). By generating target potentials in a forward manner, the mix of information encoded in each layer is expected to transition incrementally from predominantly input information in early layers to label information in later layers. We propose to define a local loss function $\mathcal{L}_{l}:=\lVert\mathbf{Z}_{l}-\tilde{\mathbf{Z}}_{l}\rVert$, which may be employed as an objective to optimise $\mathbf{W}_{l}$. Synaptic weights for each layer can be computed in a closed-form forward manner, using ridge regression (Figure 1-C), such that
$\mathbf{W}_{l}:=(\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}+\lambda\mathbf{I})^{-1}(\mathbf{A}_{l-1}^{\top}\tilde{\mathbf{Z}}_{l}).$ (2)
@@
Here, $\mathbf{A}_{l-1}\in\mathbb{R}^{N\times m_{l-1}}$ and $\tilde{\mathbf{Z}}_{l}\in\mathbb{R}^{N\times m_{l}}$ are matrices of pre-synaptic activities and target potentials, respectively, collected over the $N$ samples in $\mathcal{D}$. $\lambda$ is a regularisation term and $\mathbf{I}$ is the identity matrix. Observe that the $\mathbf{A}_{l-1}^{\top}\mathbf{A}_{l-1}$ and $\mathbf{A}_{l-1}^{\top}\tilde{\mathbf{Z}}_{l}$ terms in (2) may be computed sequentially over instances $\{\mathbf{a}_{1,l-1},\ldots,\mathbf{a}_{N,l-1}\}\subset\mathbb{R}^{1\times m_{l-1}}$, and $\{\tilde{\mathbf{z}}_{1,l},\ldots,\tilde{\mathbf{z}}_{N,l}\}\subset\mathbb{R}^{1\times m_{l}}$ (see remark A.6 in SI). Therefore, memory requirements are independent of $N$, depending only on $m_{l-1}$ and $m_{l}$. Note, weights are computed once only, after the single training epoch. The approach of encoding information on both inputs and labels in the hidden layers is inspired by the Predictive Coding and Target Propagation paradigms [5, 1, 9]. However, unlike previous approaches, no backward communication is required from neuronal outputs to pre-synaptic inputs to achieve the FP fit. As a consequence of this fitting approach, neural membrane potentials in hidden layers of FP-trained models may be interpreted as label predictions (Figure 1-C).
@@
Figure 1: Graphical overview of forward projection. A: Forward projection algorithm for fitting layer weights $\mathbf{W}_{1},\ldots,\mathbf{W}_{l}$ to model labels $\mathbf{Y}$ from data $\mathbf{X}$. B: Procedure for generating the $l$-th layer target potentials $\tilde{\mathbf{Z}}_{l}$. Pre-synaptic inputs $\mathbf{A}_{l-1}$ and labels $\mathbf{Y}$ are projected with fixed matrices $\mathbf{Q}_{l}$ and $\mathbf{U}_{l}$, respectively, before applying non-linearity $g_{l}$. C: Optimising $\mathbf{W}_{l}$ to predict $\tilde{\mathbf{Z}}_{l}$ from $\mathbf{A}_{l-1}$ by ridge regression with penalty $\lambda$. D: Interpreting membrane potentials $\mathbf{z}_{l}$ as a local label prediction $\hat{\mathbf{y}}_{l}$ given pre-synaptic inputs $\mathbf{a}_{l-1}$ and projection matrices $\mathbf{Q}_{l}$ and $\mathbf{U}_{l}^{+}$, where $\mathbf{U}_{l}^{+}$ is the pseudo-inverse of $\mathbf{U}_{l}$.